Refrigerant 134a enters a well-insulated nozzle at 14 bar, 60°C, with a velocity of 40 m/s and exits at 1.2 bar with a velocity of 460 m/s. For steady-state operation, and neglecting potential energy effects, determine: (a) the exit temperature, in K.
Answer #1

From Steady state energy equation,

Q = Wcv+ change in K.E + change in P.E + change in Enthalpy

As it is given that the nozzle is well insulated and we have to neglect P.E effect

we have,

change in K.E + change in Enthapy = 0

1/2[m(v22 - v12)] = - m(h2-h1)

1/2(v22 - v12) = (h1-h2) = cp(t1-t2)

given : v1=40 m/s , v2= 460m/s , t1= 273+60=333 K

From psychomateric tables , At P=14 bar and temp = 600C , Isobaric Specific Enthalpy = 1.2092 kJ/kg-K

1/2(4602-402) = 1.2092*1000(333-t2)

t2 = 246.165 K